OpenStudy (ineedhelpnow):

the average salary for computer programmers is $7740 less than twice the average salary for carpenters. combined their average salaries are $99,000. determine the average salary for each of these jobs.

OpenStudy (mathteacher1729):

You need to set up two equations in two variables here. Question (1): What should your variables be?

OpenStudy (ineedhelpnow):

not sure. help?

OpenStudy (mathteacher1729):

Well, you have two salaries... Whose salaries are they?

OpenStudy (ineedhelpnow):

computer programmers and carpenters

OpenStudy (anonymous):

Let x be computer programmers and y be carpenters x = 1/2y - 7740 x + y = 99000 (1/2y - 7740) + y =99000 3/2y = 91260 3y = 182520 y = 60840 x + (60840) = 99000 x = 38160